# Scalable learning of macroscopic stochastic dynamics

### v1

@@ 3.3 Ising model
$|dU/dL|_{T=T_{c}}\sim L^{1/\nu}.$ (27)
We plot the raw Binder cumulant data in Fig.˜6 (c) and compute the slope at $T_{c}$ by fitting a curve to the data. In Fig.˜6 (d), a log-log fit gives the estimated $\nu^{\ast}=0.95$, which is close to the theoretical value of $\nu=1$.
From Eq.˜23, we note that if we plot $\langle|M|\rangle L^{\beta/\nu}$ as a function of $(T-T_{c})L^{1/\nu}$, the data should collapse onto the same curve for different $T$ and $L$. Similarly, $\mathcal{X}L^{-\gamma/\nu}$ and $(T-T_{c})L^{1/\nu}$ should also collapse onto the same curve. We present these results in Fig.˜7. We can observe that even though the raw magnetization data and raw susceptibility data lie on different curves for different $T$ and $L$, the scaled data approximately collapse onto the same curve. This validates that our method can accurately capture the critical behavior of macroscopic dynamics in the Ising model.
3.4 NbMoTa alloy
@@
The microscopic diffusion dynamics are modeled using the Kinetic Monte Carlo (KMC) algorithm. In body-centered cubic (BCC) systems such as NbMoTa alloy, each vacancy $i$ can jump to one of its eight first-nearest neighbors with rate $k_{ij}=k_{0}\exp(-E_{ij}/k_{B}T),j=1,\cdots 8$, where $E_{ij}$ is the energy barrier for the jump to the $j$-th neighbor and $k_{0}$ is an attempt frequency. In our study, we will employ the pretrained neural network model from Ref. [39] for energy barrier calculation. At each KMC step, the transition rates for all vacancies are computed, giving the total jump rate $R_{\text{tot}}=\sum_{i,j}k_{ij}$. Next, the time increment is sampled from the exponential distribution $\Delta t\sim-\frac{\ln r}{R_{\text{tot}}},r\sim\mathcal{U}(0,1)$. For large systems, the number of vacancies is substantial, and evaluating transition rates for all possible events is computationally demanding.
@@
In our experiment, we set the small-system size to $n_{s}=1,024$, with a vacancy concentration corresponding to $1$ vacancy per $1024$ sites. Denote the supercell length by $L$, then the supercell will consist of $2L^{3}$ atoms for a BCC lattice. Hence, the supercell length of the small system is $L=8$. Within the hierarchical upsampling scheme, an Upsample step is performed by concatenating multiple small-system configurations, followed by a LocalRelax step where KMC dynamics are run within small patches to remove unphysical artifacts. We first train an SDE model for the small system with $n_{s}=1024$ atoms, and find that learning the macroscopic dynamics of the $6$ macroscopic observables directly can already give accurate results consistent with microscopic simulations. Therefore, for the macroscopic dynamics identification of large system from data $\mathcal{D}$, no closure variables are identified and the latent dimension is $6$.
@@
The trained SDE is subsequently simulated for a long time at different temperatures. Fig.˜8 (b) shows that the absolute value of $\delta_{\text{Mo-Ta}},\delta_{\text{Ta-Ta}},\delta_{\text{Mo-Mo}}$ increases for $T<800K$, reach a maximum around $T=800\text{-}900K$, and decrease thereafter. The results indicate a critical temperature around $800\sim 900K$, corresponding to the regime of maximal diffusion-favored ordering. Our results correspond to the results obtained from the microscopic simulation in Ref. [39], confirming the accuracy and effectiveness of our method in detecting phase transitions.
@@
To assess scalability, we further extend our method to much larger systems with $n=65,536(L=32)$ and $n=524,288(L=64)$ atoms at $T=2000K$. In Fig.˜8 (c), we show the macroscopic dynamics of $\delta_{\text{Ta-Ta}}$ for different system sizes at $T=2000K$. While larger systems require more KMC time steps to reach equilibrium, each KMC step corresponds to a smaller time increment. From Fig.˜8(c), we observe that the equilibrium KMC time is nearly identical across systems of different sizes. While the mean of the macroscopic dynamics converges approximately to the same value, the trajectories of smaller systems are more stochastic.
4 SUMMARY AND OUTLOOK
@@
This work proposed a framework to learn macroscopic dynamics of large microscopic systems from small-system simulations. We apply our method to SPDEs, spin systems, and an NbMoTa alloy system, scaling it to a large system with $524,288$ atoms. Through these applications, we highlight the capability of our model to capture accurate macroscopic dynamics over a wide range of temperatures and system sizes.
@@
Looking ahead, this framework holds significant promise for bridging the gap between microscopic simulations and macroscopic material behavior for complex material systems, such as high-entropy alloys [43, 44, 45] and polymer solutions [46, 47]. However, when applied to extremely large systems with billions of atoms, the hierarchical upsampling scheme requires many iterations, and the quality of the generated data distribution $\mathcal{D}$ may deteriorate as the number of iterations increases. We will explore more efficient strategies for generating large-system data distributions $\mathcal{D}$ to mitigate this challenge. By doing so, we envision our framework serving as a powerful tool to accelerate the discovery and design of advanced functional materials for applications in energy storage, catalysis, and structural technologies.
Code Availability
Acknowledgments
@@
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028), the Ministry of Education, Singapore, under its funding for the Research Centre of Excellence Institute for Functional Intelligent Materials (Project No. EDUNC-33-18-279-V12), and Academic Research Fund Tier 3 Grant (Project No. MOET32024-0002).
Appendix
Appendix A Continous-time Glauber dynamics
@@
Appendix B Theoretical Analysis
Proof of Theorem˜1.
@@ Appendix B Theoretical Analysis
$\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\,\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2\log p\big(\mathbf{z}_{t+\delta t}\,\big|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t\big)\big],$
$\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\,\mathbb{E}_{\mathcal{I}}\,\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[-2\log p\big(\mathbf{z}_{t+\delta t,\mathcal{I}}\,\big|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,K\,\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t\big)\big].$
Denote $\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}-\boldsymbol{\mu}(\mathbf{z}_{t})\delta t$ by $\boldsymbol{\delta}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t})$ and $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}-\boldsymbol{\mu}(\mathbf{z}_{t})\delta t$ by $\boldsymbol{\delta}_{\mathcal{I}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}})$. Since $p$ represents the probability distribution of a multivariate normal distribution, we have:
$\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log|\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t|\big]+\boldsymbol{\delta}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t})^{T}(\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t)^{-1}\boldsymbol{\delta}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t}),$
$\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log|K\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t|+\boldsymbol{\delta}_{\mathcal{I}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}})^{T}(K\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t)^{-1}\boldsymbol{\delta}_{\mathcal{I}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}})\big].$
For further simplicity, replace $\boldsymbol{\Sigma}$ by $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ and omit the dependency of $\boldsymbol{\delta},\boldsymbol{\delta}_{\mathcal{I}},\boldsymbol{\mu},\Lambda$ on $\mathbf{z}_{t},\mathbf{z}_{t+\delta t},\mathbf{z}_{t+\delta t,\mathcal{I}}$, then:
$\quad\hat{\mathcal{L}}[\boldsymbol{\mu},\boldsymbol{\Lambda}]=\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log(\delta t)-\log|\boldsymbol{\Lambda}|+\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\boldsymbol{\delta}\big],$
$\quad\hat{\mathcal{L}}_{p}[\boldsymbol{\mu},\boldsymbol{\Lambda}]=\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log(K\delta t)-\log|\boldsymbol{\Lambda}|+\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\boldsymbol{\delta}_{\mathcal{I}}\big].$
We will proceed to calculate the first-order variation and second-order variation of $\hat{\mathcal{L}}$:
@@
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log(\delta t)-\log|\boldsymbol{\Lambda}+\epsilon\mathbf{H}|+\tfrac{1}{\delta t}(\boldsymbol{\delta}-\epsilon\mathbf{h}\delta t)^{T}(\boldsymbol{\Lambda}+\epsilon\mathbf{H})(\boldsymbol{\delta}-\epsilon\mathbf{h}\delta t)\big]$
@@
$\quad+\epsilon\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]+\tfrac{\epsilon^{2}}{2}\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-4\boldsymbol{\delta}^{T}\mathbf{H}\mathbf{h}\big]+o(\epsilon^{2}).$
@@
$\log|\boldsymbol{\Lambda}+\epsilon\mathbf{H}|$ $=\log|\boldsymbol{\Lambda}|+\epsilon\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})-\tfrac{\epsilon^{2}}{2}\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+o(\epsilon^{2}),$
we have:
@@
$=\hat{\mathcal{L}}[\boldsymbol{\mu},\boldsymbol{\Lambda}]+\epsilon\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]$
$+\tfrac{\epsilon^{2}}{2}\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-4\boldsymbol{\delta}^{T}\mathbf{H}\mathbf{h}\big]+o(\epsilon^{2}).$

### v2

@@ III.2 Ising model
$|dU/dT|_{T=T_{c}}\sim L^{1/\nu}.$ (29)
We plot the raw Binder cumulant data in Fig. 5 (c) and compute the slope at $T_{c}$ by fitting a curve to the data. In Fig. 5 (d), a log-log fit gives the estimated $\nu^{\ast}=0.97$, which is close to the theoretical value of $\nu=1$.
From Eq. 25, we note that if we plot $\langle|M|\rangle L^{\beta/\nu}$ as a function of $(T-T_{c})L^{1/\nu}$, the data should collapse onto the same curve for different $T$ and $L$. Similarly, $\mathcal{X}L^{-\gamma/\nu}$ and $(T-T_{c})L^{1/\nu}$ should also collapse onto the same curve. We present these results in Fig. 6. We can observe that even though the raw magnetization data and raw susceptibility data lie on different curves for different $T$ and $L$, the scaled data approximately collapse onto the same curve. This validates that our method can accurately capture the critical behavior of macroscopic dynamics in the Ising model.
III.3 NbMoTa alloy
@@
The microscopic diffusion dynamics are modeled using the Kinetic Monte Carlo (KMC) algorithm. In body-centered cubic (BCC) systems such as NbMoTa alloy, each vacancy $i$ can jump to one of its eight first-nearest neighbors with rate $k_{ij}=k_{0}\exp(-E_{ij}/k_{B}T),j=1,\cdots 8$, where $E_{ij}$ is the energy barrier for the jump to the $j$-th neighbor and $k_{0}$ is an attempt frequency. In our study, we employ the pretrained neural network model from Ref. [56] for energy barrier calculation. At each KMC step, the transition rates for all vacancies are computed, giving the total jump rate $R_{\text{tot}}=\sum_{i,j}k_{ij}$. Next, the time increment is sampled from the exponential distribution $\Delta t\sim-\frac{\ln r}{R_{\text{tot}}},r\sim\mathcal{U}(0,1)$. For large systems, the number of vacancies is substantial, and evaluating transition rates for all possible events is computationally demanding.
@@
In our experiment, we set the small-system size to $n_{s}=1,024$, with a vacancy concentration corresponding to $1$ vacancy per $1024$ sites. Denote the supercell length by $L$, then the supercell will consist of $2L^{3}$ atoms for a BCC lattice. Hence, the supercell length of the small system is $L=8$. Within the hierarchical upsampling scheme, an Upsample step is performed by concatenating multiple small-system configurations, followed by a LocalRelax step where KMC dynamics are run within small patches to remove unphysical artifacts. We first train an SDE model for the small system with $n_{s}=1024$ atoms, and find that learning the macroscopic dynamics of the $6$ macroscopic observables directly can already give accurate results consistent with microscopic simulations. Therefore, for the macroscopic dynamics derivation of the large system from data $D$, no closure variables are derived and the latent dimension is $6$.
@@
The trained SDE is subsequently simulated for a long time at different temperatures. Fig. 7 (b) shows that the absolute value of $\delta_{\text{Mo-Ta}},\delta_{\text{Ta-Ta}},\delta_{\text{Mo-Mo}}$ increases for $T<800K$, reaches a maximum around $T=800\text{-}900K$, and decreases thereafter. The results indicate a critical temperature around $800\sim 900K$, corresponding to the regime of maximal diffusion-favored ordering. Our results correspond to the results obtained from the microscopic simulation in Ref. [56], confirming the accuracy and effectiveness of our method in detecting phase transitions.
@@
To assess scalability, we further extend our method to much larger systems with $n=65,536(L=32)$ and $n=524,288(L=64)$ atoms at $T=2000K$. In Fig. 7 (c), we show the macroscopic dynamics of $\delta_{\text{Ta-Ta}}$ for different system sizes at $T=2000K$. While larger systems require more KMC time steps to reach equilibrium, each KMC step corresponds to a smaller time increment. From Fig. 7(c), we observe that the equilibrium KMC time is nearly identical across systems of different sizes. While the mean of the macroscopic dynamics converges approximately to the same value, the trajectories of smaller systems are more stochastic.
IV DISCUSSION
@@
This work proposes a framework to learn macroscopic dynamics of large microscopic systems from small-system simulations. We apply our method to SPDEs, spin systems, and an NbMoTa alloy system, scaling it to a large system with 524,288524{,}288 atoms. Through these applications, we highlight the capability of our model to capture accurate macroscopic dynamics over a wide range of temperatures and system sizes.
The conventional multiscale workflow typically begins by assuming a specific continuum evolution equation a priori and then determines the associated coefficients from experiments or atomistic simulations. Common choices for the continuum evolution equation include Fick’s laws [17], Onsager transport equations [18], and phase-field models such as the Allen-Cahn and Cahn-Hilliard equations [8, 2, 10]. For example, diffusion is typically described by Fick’s law $\partial_{t}c=\nabla\!\cdot\!\big(D\nabla c\big)$, where the diffusion coefficient is subsequently estimated from the mean squared displacement of molecular dynamics trajectories [52, 42].
Both our workflow and the conventional multiscale workflow aim to construct effective dynamical descriptions from microscopic simulations to reduce the computational cost of tracking massive atomic degrees of freedom. However, our framework differs in several key respects. First, conventional multiscale approaches are typically formulated at a mesoscopic level, while our method learns the macroscopic dynamics. In conventional multiscale approaches, the spatially resolved field variables depend locally on microscopic coordinates. In contrast, our method explicitly targets the macroscopic dynamics. In our framework, the observables of interest depend globally on all the microscopic coordinates, resulting in macroscopic dynamics that evolve solely as a function of time rather than space. While we utilize an encoder $\boldsymbol{\varphi}$ to extract latent features $\boldsymbol{\varphi}(\mathbf{x}_{t,\mathcal{I}})$ from each patch, these features are aggregated to define the state $\mathbf{z}_{t}=\boldsymbol{\varphi}(\mathbf{x}_{t})$ for the whole large system. Our method explicitly models the dynamics of $\mathbf{z}_{t}$ from short-time, small-scale microscopic information, and is therefore best viewed as a macroscale modeling approach. Second, regarding the model structure, conventional multiscale approaches explicitly assume a predefined functional continuum evolution equation, which often limits their accuracy and ability to capture complex or non-linear phenomena. In contrast, our framework learns both the closure variables and the macroscopic dynamics directly from data, without prescribing a specific functional form. Although we employ structured architectures such as OnsagerNet, these structures serve only to enforce fundamental physical constraints, such as thermodynamic consistency, rather than to constrain the admissible functional forms of the dynamics. By parameterizing the encoder and macroscopic dynamics with neural networks, we ensure high model expressivity. Consequently, our framework offers significantly greater flexibility in discovering the underlying governing laws directly from data.
In our framework, we enforce physical consistency of the macroscopic dynamics by modeling the SDE using a stochastic OnsagerNet. The closure variables, on the other hand, are obtained through data-driven learning and may contain redundancy or lack direct physical interpretability. To reduce redundancy, one practical approach is to vary the dimension of the closure variables and identify when further increases no longer improve macroscopic prediction accuracy. In addition, physical interpretability may be improved by incorporating basic physical or symmetry constraints into the encoder, such as translation invariance for lattice systems with periodic boundary conditions, which we leave for future investigation.
A practical limitation of the present framework arises from the hierarchical upsampling procedure used to generate large-system snapshots from small-system simulations. The LocalRelax step is used to remove the unphysical artifacts introduced by the Upsample step. In this paper, LocalRelax is implemented by dividing each upsampled large-system snapshot into overlapping patches and applying short-time local dynamics evolution within each patch. This simple procedure is effective in our experiments for removing common unphysical artifacts introduced by upsampling. However, for more complex systems, this simple LocalRelax step may be insufficient to adequately remove the unphysical artifacts. Moreover, when applied to extremely large systems with billions of atoms, the hierarchical upsampling scheme requires many iterations, and the quality of the generated dataset $D$ may deteriorate as the number of iterations increases. The hierarchical upsampling strategy adopted in this work represents only one possible approach, and more sophisticated strategies may be required for more complex systems. We acknowledge this as a limitation of the current method. In the future, we will explore more efficient strategies for generating large-system datasets $D$ to mitigate this challenge.
@@
Looking ahead, this framework holds significant promise for bridging the gap between microscopic simulations and macroscopic material behavior in complex material systems, such as high-entropy alloys [13, 38, 9] and polymer solutions [39, 27]. We envision this framework as a powerful tool for accelerating the discovery and design of advanced functional materials, with potential applications in energy storage, catalysis, and structural technologies.
Acknowledgements.
@@
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028), the Ministry of Education, Singapore, under its funding for the Research Centre of Excellence Institute for Functional Intelligent Materials (Project No. EDUNC-33-18-279-V12), and Academic Research Fund Tier 3 Grant (Project No. MOET32024-0002). The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg).
Data Availability
The supporting data and code are available in the public repository [1].
Appendix A Notation and Symbols
The notation and symbols are summarized in Table 1.
Symbol Meaning
$\mathbf{x}_{t}$ Microscopic state of the large system.
$\mathcal{I}$ Index set of a randomly selected patch within the large system.
$\delta t$ Time step.
$x_{t,\mathcal{I}}$ Microscopic state restricted to patch $\mathcal{I}$ at time $t$.
$x_{t+\delta t,\mathcal{I}}$ Microscopic state restricted to patch $\mathcal{I}$ at time $t+\delta t$.
$n_{s}$ Number of lattice sites in the small system.
$n$ Number of lattice sites in the large system.
$\mathcal{S}_{n_{s}}$ Microscopic simulator that can accurately simulate the microscopic dynamics of a small system up to $n_{s}$ lattice sites.
$D_{s}$ Dataset of small-system snapshots.
$D$ Training dataset constructed from $D_{s}$ through hierarchical upsampling scheme.
$K$ Number of patches used to partition the large system.
$\boldsymbol{\varphi}^{\ast}$ Mapping from a microscopic state to macroscopic observables.
$\hat{\boldsymbol{\varphi}}$ Mapping from a microscopic state to closure variables.
[…]
$\boldsymbol{\psi}$ Decoder.
$\boldsymbol{\mu}$ Drift term of the latent dynamics.
$\boldsymbol{\Sigma}$ Diffusion term of the latent dynamics.
$\mathbf{z}^{\ast}$ Macroscopic observables defined by $\mathbf{z}^{\ast}=\boldsymbol{\varphi}^{\ast}(\mathbf{x})$
$\hat{\mathbf{z}}$ Closure variables defined by $\hat{\mathbf{z}}=\hat{\boldsymbol{\varphi}}(\mathbf{x})$
$\mathbf{z}_{t}$ Latent state defined by $\mathbf{z}_{t}=\boldsymbol{\varphi}(\mathbf{x}_{t})$.
$\mathbf{z}_{t+\delta t,\mathcal{I}}$ Latent state corresponding to $x_{t+\delta t,\mathcal{I}}$.
$\Delta\mathbf{z}_{t}$ Increment of the latent state over $\delta t$, defined by $\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}$
$\Delta\mathbf{z}^{\ast}_{t}$ Increment of the macroscopic observable, defined by $\mathbf{z}^{\ast}_{t+\delta t}-\mathbf{z}^{\ast}_{t}$
$\Delta\mathbf{z}_{t,\mathcal{I}}$ Increment of the latent state restricted to patch $\mathcal{I}$, defined by $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t,\mathcal{I}}$
$\Delta\mathbf{z}^{\ast}_{t,\mathcal{I}}$ Increment of the macroscopic observable restricted to patch $\mathcal{I}$, defined by $\mathbf{z}^{\ast}_{t+\delta t,\mathcal{I}}-\mathbf{z}^{\ast}_{t,\mathcal{I}}$
Table 1: Summary of key symbols and definitions.
Appendix B Continous-time Glauber dynamics
The algorithm of continuous-time Glauber dynamics is summarized in Algorithm 2.
@@ Appendix B Continous-time Glauber dynamics
Appendix C Theoretical Analysis
Proof of Theorem 1.
We divide the proof into two parts: first, we establish the closed-form expressions for the unique minimizers using variational calculus; second, we derive the error bounds based on the total variation distance.
1. Derivation of Minimizers
@@
$\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\,\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[$
$-2\log p\big(\mathbf{z}_{t+\delta t}\,\big|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t\big)\big],$
$\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\,\mathbb{E}_{\mathcal{I}}\,\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[$
$-2\log p\big(\mathbf{z}_{t+\delta t,\mathcal{I}}\,\big|\,\mathbf{z}_{t}+\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,\,K\,\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t\big)\big].$
Recall $\Delta\mathbf{z}_{t}=\mathbf{z}_{t+\delta t}-\mathbf{z}_{t},\Delta\mathbf{z}_{t,\mathcal{I}}=\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}$. By introducing the precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ and the displacement vector
$\boldsymbol{\delta}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t})\coloneqq\Delta\mathbf{z}_{t}-\boldsymbol{\mu}(\mathbf{z}_{t})\delta t$
$\boldsymbol{\delta}_{\mathcal{I}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}})\coloneqq\Delta\mathbf{z}_{t,\mathcal{I}}-\boldsymbol{\mu}(\mathbf{z}_{t})\delta t,$
we can rewrite the objective as:
$\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log|\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t|\big]$
$+\boldsymbol{\delta}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t})^{T}(\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t)^{-1}\boldsymbol{\delta}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t}),$
$\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[d\log(2\pi)$
$+\log|K\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t|$
$+\boldsymbol{\delta}_{\mathcal{I}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}})^{T}(K\boldsymbol{\Sigma}(\mathbf{z}_{t})\delta t)^{-1}\boldsymbol{\delta}_{\mathcal{I}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t,\mathcal{I}})\big].$
For further simplicity, we omit the notational dependency of $\boldsymbol{\delta},\boldsymbol{\delta}_{\mathcal{I}},\boldsymbol{\mu},\Lambda$ on $\mathbf{z}_{t},\mathbf{z}_{t+\delta t},\mathbf{z}_{t+\delta t,\mathcal{I}}$. We will proceed to calculate the first-order variation and second-order variation of $\hat{\mathcal{L}}$:
@@
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[d\log(2\pi)+\log(\delta t)-\log|\boldsymbol{\Lambda}+\epsilon\mathbf{H}|$
$\quad+\tfrac{1}{\delta t}(\boldsymbol{\delta}-\epsilon\mathbf{h}\delta t)^{T}(\boldsymbol{\Lambda}+\epsilon\mathbf{H})(\boldsymbol{\delta}-\epsilon\mathbf{h}\delta t)\big]$
@@
$\quad+\epsilon\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]$
$\quad+\tfrac{\epsilon^{2}}{2}\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-4\boldsymbol{\delta}^{T}\mathbf{H}\mathbf{h}\big]+o(\epsilon^{2}).$
@@
$\log|\boldsymbol{\Lambda}+\epsilon\mathbf{H}|$ $=\log|\boldsymbol{\Lambda}|+\epsilon\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$-\tfrac{\epsilon^{2}}{2}\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+o(\epsilon^{2}),$
@@
$=\hat{\mathcal{L}}[\boldsymbol{\mu},\boldsymbol{\Lambda}]$
$+\epsilon\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]$
$+\tfrac{\epsilon^{2}}{2}\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$\quad+2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-4\boldsymbol{\delta}^{T}\mathbf{H}\mathbf{h}\big]$
$+o(\epsilon^{2}).$